OpenStudy (anonymous):

express sin45 in term of sin45 and cos45,evaluate sin46 and estimate your error

OpenStudy (anonymous):

good morning my people i need answer to that question. thanks

OpenStudy (callisto):

sin45 = 1/2 ( sin45 +cos45) From calculator , sin46 = 0.719, cor to 3 d.p error of my answer = +/-0.001/2 = +/-0.0005

OpenStudy (callisto):

Hmm.. sin45 = cos 45 So , sin45 = 1/2 ( sin45 +cos45)

OpenStudy (anonymous):

@callisto,work it out in term ot maclaurin

OpenStudy (experimentx):

you mean maclaurin series??

OpenStudy (anonymous):

oh... maclaurin...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@order i mean maclaurin series

OpenStudy (experimentx):

sin x = x - x^3/3! + x^5/5! - x^7/7! + .... sin (pi/4) = (pi/4) - (pi/4)^3/3! + (pi/4)^5/5! - (pi/4)^7/7! + (pi/4)^9/9! = 0.7071 ... says wolf http://www.wolframalpha.com/input/?i=%28pi%2F4%29+-+%28pi%2F4%29%5E3%2F3%21+%2B+%28pi%2F4%29%5E5%2F5%21+-+%28pi%2F4%29%5E7%2F7%21+%2B+%28pi%2F4%29%5E9%2F9%21&a=ClashPrefs_*MathWorld.MaclaurinSeries-

OpenStudy (anonymous):

@experimentx use this formula f(x)=f(a)+hf'(a)+h^2/2!f"(a)

OpenStudy (experimentx):

oo ... that quadratic approximation.

OpenStudy (experimentx):

and what points are we supposed to use inplace of a is it 0?? i mean for 45??

OpenStudy (anonymous):

sin45^0

OpenStudy (experimentx):

sin(x)=sin(a)+x cos(a) - x^2/2! sin(a) sin(pi/4)=sin(0)+pi/4 cos(0) - x^2/2! sin(0) = pi/4

OpenStudy (anonymous):

no

OpenStudy (experimentx):

so ??

OpenStudy (experimentx):

what point are we going to use as a ... we need a reference point right? i look the liberty of taking a=0